Write the slope intercept equation for the line that passes through (14,-2) and is perpendicular to 7x-10y=18
Answer by alicealc(293)

7x - 10y = 18
7x - 18 = 10y
(divide all by 10)
7/10 x - 18/10 = y
the slope for this line is 7/10

2 lines are perpendicular if
slope line 1 * slope line 2 = -1
7/10 * slope line 2 = -1
slope line 2 = -1 * 10/7 = -10/7

line equation that has slope of m and passes through (x1, y1):
y - y1 = m*(x - x1)

this line has the slope of -10/7 and passes through (14, -2), so:
y - (-2) = -10/7 * (x - 14)
y + 2 = -10/7 x + 140/7
y + 2 = -10/7 x + 20
y = -10/7 x + 20 - 2
y = -10/7 x + 18